OpenStudy (anonymous):

Factorising the following expression: X^2 + 8x + 4 - y^2 Which method of factorising is the most appropriate? Cheers

OpenStudy (mertsj):

Not factorable

OpenStudy (anonymous):

I think the "completing the square" method is your best option. So: \[=x ^{2}+8x+(8\div2)^{2}-(8\div2)^{2}-y^{2}\] \[=(x+4)^{2}-4^{2}-y ^{2}+4\] \[=(x+4)^{2}-y^{2}-4^{2}+4\] \[=(x+4-y)(x+4+y)-12\] And I think that's it. There's not much else you can do.

OpenStudy (mertsj):

Bit that is not factored!!

OpenStudy (mertsj):

We may as well write x(x+8)-(y-2)(y+2)

OpenStudy (anonymous):

Well it depends on the problem Chad123 is trying to solve. I understand it is not "fully" factorized, but perhaps it will help them anyway.

OpenStudy (anonymous):

Thanks guys for your help. I'll give each one of you a medal :)